assume that you place money in an account that is compounded annually. how long will it take the investment to increase by 60% at an interest rate of 5%
Answer by lwsshak3(11628)

assume that you place money in an account that is compounded annually. how long will it take the investment to increase by 60% at an interest rate of 5%
***
Compound Interest Formula: A=P(1-i)^n, P=initial investment, i=interest per period, n=number of periods, A=amount after n-periods
For given problem:
i=5%
A/P=1.6=(1+i)^n
1.6=(1+.05)^n
take log of both sides:
log(1.6)=nlog(1.05)
n=log(1.6)/log(1.05)
n≈9.6 yrs